MelSaverethiar:

1 5 sinh x = 1 9 sinh y = 1 28 sinh(x + y),

karissafrazier:

huh?

MelSaverethiar:

Solve the equations 1/5 sinh x = 1/9 sinh y = 1/28sinh (x + y), giving your answers to 3 d.p. please

karissafrazier:

i can try my best

MelSaverethiar:

yess ty

MelSaverethiar:

im not understanding. but thank u for ur time

MelSaverethiar:

i m in university 1st year

CheeseStick7793:

x=\ln \left(\frac{1+\sqrt{226}}{15}\right)

CheeseStick7793:

CheeseStick7793:

here is your answer

MelSaverethiar:

thank you so much everyone thx @cheesestick7793

CheeseStick7793:

mhm